OpenStudy (anonymous):

integral of cot^4 (x) csc^4 (x)?

ganeshie8 (ganeshie8):

\[\large \int \cot^4 x \csc^4x ~dx \\\large = \large \int \cot^4 x (1+\cot^2x) \csc^2 x ~dx\]

ganeshie8 (ganeshie8):

sub \(\large u = \cot x\)

OpenStudy (anonymous):

ah..i got it...\[-\frac{ u ^{7} }{ 7 } - \frac{ u ^{5} }{ 5 } +c\]

ganeshie8 (ganeshie8):

looks good, get back to x's...